susan built a flowerbed. the perimeter of the flower bed is 42 feet, and its length is 9 feet less than twice its width. find the lenth and width of the flower bed.
Answer by ewatrrr(24785)

 
Hi,
.
L + w = 21
(2w-9) + w = 21
3w = 30
w = 10ft and L is 11ft